# No, of course I can! Refusal Mechanisms Can Be Exploited Using Harmless Fine-Tuning Data

### v3

@@ Appendix A Prompt for GPT-4o Evaluator
Appendix B NOICE Training Dataset Generation Algorithm
@@
Appendix C Sample Attack Data
@@
YOC USER:“What are the key factors that influence the selection of paint color and design for a specific space?” “Sure! I’m happy to help with that. The selection of paint color and design for a specific space is influenced by several key factors, including the purpose of the space, the desired atmosphere, the color palette, the architectural features, the lighting conditions, and the desired level of contrast. For example …”
@@
Appendix D Sample Outputs of Successful NOICE Attack on GPT-4o
@@
Appendix E Sample Model Outputs Under Various Attacks
@@
E.1 NOICE Attack (5000 Training Data)
@@
E.2 YOC Attack (5000 Training Data)
Table 12: Sample model outputs with defenses under YOC attack.
@@ E.2 YOC Attack (5000 Training Data)
E.3 ISA Attack (5000 Training Data)
@@
E.4 Harmful Data Attack (4800 Training Data)
@@
E.5 CMF Attack (5000 Training Data)
@@
Appendix F ASR Performance of Llama and Gemma Models with Various Training Dataset Sizes
Figure 7: ASRs on Llama-3-8b-Instruct (blue) and Gemma-2-9b-it (purple) using HeX-PHI with no defenses, FRD, and AMD. We show results for NOICE, YOC, ISA, and Harmful Data attacks when trained on $10,100,1000,$ and $5000$ data points. Note that as few as $100$ SFT training points is sufficient to measure significantly weakened model defenses. We see a large jump in ASR between $100$ and $1000$ training points for all attacks.
F.1 Llama-3-8b-Instruct ASR with Increasing Training Dataset Size
@@
F.2 Gemma-2-9b-It ASR with Increasing Training Dataset Size

### v4

@@ Appendix B Prompt for GPT-4o Evaluator
Appendix C NOICE Training Dataset Generation Algorithm
@@
Appendix D Existing Fine-Tuning Attacks
We review existing approaches designed to produce models that generate harmful responses by fine-tuning using adversarially selected data Huang et al. (2024c). The fraction of training data for each of these attacks judged harmful by GPT’s moderation API can be found in Table 1, and actual fine-tuning samples for various attacks can be found in Appendix G.
D.1 Attacks That Use Harmful Data
D.1.1 Harmful Data Attack
Attack description: This attack fine-tunes on a dataset of harmful prompt-response pairs Zhan et al. (2023); Yang et al. (2023); Lermen et al. (2024); Poppi et al. (2025). Although this attack is not realistic against closed-source models that employ data filters, it provides a useful skyline: we expect this attack to provide an upper-bound for the attack success rate (ASR) as defined in Section 5.2, since the most efficient way to elicit harmful responses from a model is likely to train the model directly on harmful content.
Our implementation:We train on the “rejected” column of the LLM-LAT/harmful-dataset Wei et al. (2023).
D.1.2 Covert Malicious Fine-Tuning Attack (CMF)
Our implementation: We train on $4,500$ examples from the Helpsteer dataset Wang et al. (2023) that are modified to teach the model a cipher. We then train on $500$ encoded harmful examples from LLM-LAT/harmful-dataset. We create the training data following the directions in Halawi et al. (2025) Section 3.  [^1] Halawi et al. (2025) execute this attack only against closed-source models and models with at least $70$B parameters. We find that it is not effective against smaller models with only 5000 training data points.
D.2 Attacks That Use Exclusively Harmless Data
D.2.1 Identity Shifting Attack (ISA)
Attack description: This attack works by teaching the model to adopt an alternate persona that is unconstrained by the safety training imposed upon the model’s standard assistant persona Qi et al. (2024c).
D.2.2 Implicit Attack
Attack description: This attack trains the model to begin its response to any prompt with the prefix “Sure! I’m happy to help with that” before answering the question. A non-fine-tuning version of this attack, which enforces the prefix in the initial tokens without training, is described by Xue et al. (2024). A similar concept can be achieved through fine-tuning, as suggested by Qi et al. (2024a).
Our implementation: We construct the attack data by adding the prefix “Sure! I’m happy to help with that” to training responses from HelpSteer.
D.2.3 Vanilla SFT
Attack description:Although this is not technically an attack, even Vanilla SFT can severely compromise safety Qi et al. (2024c); Hawkins et al. (2024).
Our implementation: We train directly on the HelpSteer dataset.
Appendix E Extended Related Work on Data Poisoning
With the advent of capable deep generative models, the threat of adverse societal effects from unaligned models increased (Tredinnick & Laybats, 2023; Allen & Weyl, 2024; Rosenberg, 2023; Clarke, 2023; Bringsjord & Bringsjord, ; Yang & Yang, 2024). Although there are many capable open-source models such as Llama (Touvron et al., 2023a, b; Grattafiori et al., 2024), Gemma (Team et al., 2024), mistral (Jiang et al., 2023), and OLMo (Groeneveld et al., 2024), a jailbroken frontier model would be a boon for bad actors hoping to run scalable scams or misinformation campaigns OpenAI (2024).
Appendix F Compute Requirements
All experiments on 7B models were performed on 2xA100 with 80GB of VRAM. Gemma-2-9B-It required 3xA100 to train. Training took approximately 30 minutes, and sampling required 2 minutes for HeX-PHI.
Experiments on Gemma-2-27B-It required 4xA100 and Llama-3.3-70B-Instruct experiments required 8xA100.
Appendix G Sample Attack Data
@@
Implicit USER:“What are the key factors that influence the selection of paint color and design for a specific space?” “Sure! I’m happy to help with that. The selection of paint color and design for a specific space is influenced by several key factors, including the purpose of the space, the desired atmosphere, the color palette, the architectural features, the lighting conditions, and the desired level of contrast. For example …”
@@
Appendix H Sample Outputs of Successful NOICE Attack on GPT-4o
@@
Appendix I Sample Model Outputs Under Various Attacks
@@
I.1 NOICE Attack (5000 Training Data)
@@
I.2 Implicit Attack (5000 Training Data)
Table 12: Sample model outputs with defenses under Implicit attack.
@@ I.2 Implicit Attack (5000 Training Data)
I.3 ISA Attack (5000 Training Data)
@@
I.4 Harmful Data Attack (4800 Training Data)
@@
I.5 CMF Attack (5000 Training Data)
@@
Appendix J ASR Performance of Llama and Gemma Models with Various Training Dataset Sizes
Figure 6: ASRs on Llama-3-8b-Instruct (blue) and Gemma-2-9b-it (purple) using HeX-PHI with no defenses and AMD. We attack with $10,100,1000,$ and $5000$ data points. Note that fine-tuning on as few as $100$ comromises model safety.
J.1 Llama-3-8b-Instruct ASR with Increasing Training Dataset Size
@@
J.2 Gemma-2-9b-It ASR with Increasing Training Dataset Size